OpenStudy (anonymous):

Find the value of x. x=-20\2(-5)

OpenStudy (anonymous):

Multiply the denominator first 2(-5) then divide it by -20.

OpenStudy (anonymous):

0.5

OpenStudy (michele_laino):

is your expression, like this: \[x = \frac{{ - 20}}{{2 \times \left( { - 5} \right)}}\]

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

ok! then here are more steps: \[x = \frac{{ - 20}}{{2 \times \left( { - 5} \right)}} = \frac{{ - 20}}{{ - 10}} = \frac{{20}}{{10}} = ...?\]

OpenStudy (anonymous):

2

OpenStudy (michele_laino):

that's right!